# Blockchain with proof of quantum work

### v1

@@ VII Conclusions
We have proposed and successfully tested a blockchain architecture in which proof of quantum work (PoQ) is performed on quantum computers. Our approach introduces multiple quantum-based hash generation methods, with varying levels of complexity depending on the capabilities of the quantum hardware. To address the inherent probabilistic nature of quantum hash generation and validation—an unavoidable property of any quantum hash algorithm—we developed techniques to ensure blockchain stability. We experimentally tested our proposed blockchain on four internationally distributed D-Wave annealing QPUs, demonstrating that despite architectural and parametric variations, these QPUs can be programmed to generate and validate each other’s hashes in a way that allowed stable blockchains to persist over thousands of broadcast blocks and hundreds of thousands of hash validations.
@@
For this demonstration, we used the critical dynamics of 3D and biclique spin glass at a fixed size as the computational problem for PoQ. This was motivated by its role in the first experimental demonstration of quantum supremacy in a real-world problem [1] that is not based on boson or random-circuit sampling. However, this choice does not constrain the quantum problem that can be used for PoQ. Strategies, such as randomizing problem topology, dimensionality, qubit count, and annealing time, can prevent classical learning algorithms from predicting expectation values based on large datasets [46]. By choosing a sufficiently complex ensemble of experiments and sampling it cryptographically, experimental outcomes can be made weakly correlated. Additionally, hash algorithms leveraging entanglement witnesses and shadow tomography can further enhance the quantum nature of the hashing process. We note that while supremacy is desirable for PoQ, it may not be strictly necessary as long as classical competition remains prohibitively expensive, making quantum approaches the only practicable ones. In fact, one may deliberately use smaller problem instances where the ground truth can be obtained — albeit with significant computational effort — to benchmark and calibrate QPUs.
@@
A quantum blockchain could transform existing blockchain technology by significantly mitigating the substantial electricity demands of classical proof of work (PoW) mining. Electricity costs are estimated to comprise 90% to 95% of Bitcoin’s total mining expenses [47]. In our PoQ approach, energy consumption amounts to a small fraction of quantum computation cost (which is the dominant cost in the PoQ approach). For example, energy expenses account for just 0.1% of the total computation cost on D-Wave quantum computers [50]. Replacing classical PoW with quantum PoQ could therefore reduce energy costs by a factor of 1,000. This ratio may shift as quantum computation becomes more cost-efficient over time, but substantial energy savings are expected to persist. Transition to PoQ would also relocate mining from regions with low electricity costs to countries with advanced quantum computing infrastructure. Beyond demonstrating the first practical use of quantum supremacy in a critical domain, this work proves the feasibility of quantum computing in real-world tasks and suggests broader high-impact applications with today’s technology.
@@ Acknowledgment
We thank Brian Barch, Vladimir Vargas Calderon, Rahul Deshpande, Pau Farré, Fiona Hannington, Richard Harris, Emile Hoskinson, Trevor Lanting, Daniel Lidar, Peter Love, Chris Rich, and Murray Thom for fruitful discussions and comments on the manuscript.
@@ Appendix A Cross-Validation Limitations in Boson and Random-Circuit Sampling
To operate within a quantum blockchain, QPUs must cross-validate each other’s outcomes. Unstructured distributions yielding states with close to uniform random probability, such as those obtained from boson or random-circuit sampling, are unsuitable for this purpose: the expectation values of all foreseeable k-local operators remain close to zero, making them ineffective in the construction of witnesses. The only observable with a nontrivial expectation is $\left\langle\psi\right|\rho\left|\psi\right\rangle$, where $\left|\psi\right\rangle$ is a ground-truth wave function. For this reason the estimator commonly used to establish supremacy is the cross entropy [29]. Given a distribution allowing fair and efficient sampling in the computational basis ($P_{\text{QPU}}$, typically a QPU), and a well-resolved ground-truth distribution in the same basis ($P_{\text{GT}}$, used for verification), the cross entropy is quantified as
@@
Appendix B CHSH Inequality
@@ Appendix C Weaknesses and Mitigations
Many weaknesses for PoW blockchains such as Bitcoin are known, and many mitigations have been developed [48, 7, 8, 9, 10]. We focus on several attacks that might be attempted by an adversarial agent to gain reward with incomplete work. The reward might be gained deterministically, or more often probabilistically (in expectation). Neither the attacks presented, nor the mitigations, are an exhaustive list.
@@
As described in the main text, the particular proof-of-concept experiment is premised on the notion of honest network users and is susceptible to attacks. In addition to mitigations described in the main text we explore additional procedures in this appendix. We present a set of tools to protect against no-work and partial-work attacks, but optimization of the experimental ensemble and digitalization process, essential in practice, goes beyond the scope of this publication.
@@ C.2 Partial work attacks
An attack might proceed as follows. For many nonces an attacker can estimate $\{W_{\alpha}\}$ by classical methods (e.g. by approximate simulation, or machine learning). The results are ranked by the number of (robust) threshold-satisfying bits. The best candidate might be broadcast directly. The reward from this strategy is described as in the no-work scenario, but with an enhanced (classical) work cost and enhanced probability that the bits are verified. This attack is mitigated in part by the modified block structure of Appendix C.4. In particular, spoofing enhanced experimental data (beyond the $\mathcal{N}_{\rm zeros}$ proof of work bits) is impossible without quantum supremacy [1].
@@ C.2.1 Partial Quantum Work Attack
The quantum experimental set-up also allows room for efficiencies as well, which should be understood in a full application. To minimize QPU access time (costs) a miner should optimize parameters such as number of samples and programming time. They could then more efficiently search, or filter, nonces. A miner is incentivized to prefilter nonces with cheap experiments, and then seek higher confidence (via enhanced sampling) only on candidates for broadcast. In our demonstration by contrast we fix QPU access time to 1 second in mining and validation.
@@ C.3 Multi-Block Attack
Our definitions of Chainwork mean that for any path to be viable as a strongest chain there must be strictly more verified blocks than unverified blocks. Owing to this, the criteria necessary for this attack to succeed are similar to the standard 51% attack. In effect an attacker must be able to grow the chain more quickly than alternative non-fraudulent candidates, in order to undo the sum-work damage of their fraudulent block. This requires control of a significant fraction of community compute power assuming practical chain efficiency (concentrated honest mining power, up to small fluctuations).
@@ C.5 Hyperplane distribution mitigations
Figure 7: Operation of an example blockchain of 128 block broadcasts generated in the limit of many miners using projection orthogonal to $J$ with $\mathcal{N}_{zeros}=20$, and confidence-based Chainwork with $\mathcal{N}_{\rm zeros}=1$. The most recent block mined is marked in black and the strongest chain in blue.
Suppose a per-nonce classical work less than the quantum experimental cost ($C_{H}$) allows for a witness estimator ${\hat{W}}$. One mitigation is to modify the distribution of thresholds or hyperplanes (to choose $W_{0}\sim{\hat{W}}$), or to choose hyperplanes orthogonal to ${\hat{W}}$. In either case the classical approximation is rendered uninformative on the signs expected from the quantum experiment. If the uncertainty in the quantum experiment is relatively small, we might still have sufficient accuracy in quantum experiments for robust validation. A visualization of this process is shown in Fig. 11.
In the case of our quench experiments, we know that nearest-neighbor correlation are correlated with $-J$, the approximation $C_{ij}\propto-J_{ij}$ describes accurately the weak coupling (fast anneal) limit. There are significant deviations from this estimate, but any hyperplane that is well correlated with $J$ can accumulate a reliable signal from weak estimators incorporating this approximation. To mitigate for this we might choose our hyperplanes to be orthogonal to $J$. Choosing hyperplanes (or thresholds $W_{0\alpha}$) to decrease reliability of classical estimators will typically enhance the uncertainty in quantum validation, thus reduce the efficiency and enhance delays in a blockchain. This effect is shown in Fig. 15 as a function of $\mathcal{N}_{zeros}$. In Fig. 7 we show an example of a chain resistant to simple classical inference based on local coupling strength.
@@ Appendix D Experimental Implementation Details
Figure 8: Energy scales (29) of the general-access QPUs used in experiments. $\Gamma(s)$ decreases with anneal progression $s$, ${\cal J}(s)$ increases. Transverse fields are comparable on all devices, but the problem energy scale is larger in Advantage2. Advantage QPUs can be programmed with (approximately) doubling the anneal duration to simulate an Advantage2 QPU full-energy scale anneal [1].
Figure 9: Operation of an example blockchain using four generally-accessible Advantage and Advantage2 QPUs with $\mathcal{N}_{\rm zeros}=40$ and 100 miners with data collected across more than 2 weeks, with 780 total block broadcasts and using basic +/-1 Chainwork. The color scheme matches that of Figure 3.Node placement and line-thickness conventions follow Figure 4. It is apparent in the larger number of black and gray (unconfirmed) transactions, and in the increased rate of edges spanning the spiral, that this blockchain is of lower efficiency and larger delay than that presented in Figure 4. This is consistent with larger $\mathcal{N}_{\rm zeros}$ as shown in Figure 5 and 15. A temporary decrease in efficiency was also observed mid-chain owing to one of the four QPUs going offline for an hour.
Figure 10: Operation of an example blockchain using the Advantage2 QPU and the dimerized biclique spin-glass problems with $\mathcal{N}_{\rm zeros}=20$ and 100 miners, 592 block broadcasts and using basic +/-1 Chainwork. Data was collected across 10 days. The color scheme, node placemenet and line thicknesses match the conventions of Figures 3 and 4. Data was collected across more than a week.
@@
Figure 12: Validation rates with respect to experiments on different QPUs. Validation rates are higher using a common QPU, or a common QPU architecture (Advantage (Adv) or Advantage2 prototype (Adv2)). They are slightly higher on Advantage QPUs due the smaller size of the prototype Advantage2 prototype QPU combined with the bound of 1 second of QPU access time. On each QPU the probability of an outcome 1 is measured on each QPU from $20$ independent (with randomized embeddings) programmings, for 5 $J$ realizations and total of $5120$ projections. From these 20 experiments we infer bit error rates for different QPU combinations. 3 Advantage QPUs are used, with some small variability between QPUs (not shown). (top) Hyperplane are chosen with independent and normally distributed components. (bottom) With the same hyperplanes projected into the space orthogonal to $J$ (see Section C.5).
@@
Figure 15: Statistics from 5-10 chains, each of length 512-2048 using the QPU-witness resampling methodology (see Appendices D.1 and D.3). Hyperplanes are either i.i.d normally distributed, or are restricted to be orthogonal-to-$J$ (strengthening hashes against weak classical estimators). Resistance to spoofing increases uncertainty in validation, reducing efficiency. The Chainwork uses either the basic +/-1 definition, or the confidence-weighted definition with $\mathcal{N}_{\rm max}=1$ or $2$: confidence-weighted strongest chains are more efficient. (top) Mining efficiency measures the fraction of block broadcasts entering the strongest chain in the limit of many miners. (bottom) Transactions (blocks) are uncertain until agreed upon by the majority of miners. Transaction delay with 95% confidence measures the depth required for a consolidation of 95% of mining power in the limit of many miners. Stakeholders might consider blocks buried this deep within their strongest chain as immutable with 95% confidence; higher confidence implies higher delays.
@@
We performed simulations of many chains with 100 miners modeling a mining pool and transactions, or taking the limit of many miners (where no miner mines more than one block). Examples are shown in Figs. 4, 9, 10 and 7. However, only the distribution of witness statistics impacts the structure and stability of the chain, assuming no simultaneous broadcasts, via mining events and stakeholder validation. Understanding this we were able to accelerate mining (Section D.2), and simulate experiments by resampling (Section D.3). We first describe our main statistical measures of blockchain stability in Section D.1.

### v2

@@ VII Conclusions
We have proposed and successfully tested a blockchain architecture in which proof of quantum work (PoQ) is performed on quantum computers. Our approach introduces multiple quantum-based hash generation methods, with varying levels of complexity depending on the capabilities of the quantum hardware. To address the inherent probabilistic nature of quantum hash generation and validation—an unavoidable property of any quantum hash algorithm we developed techniques to ensure blockchain stability. We experimentally tested our proposed blockchain on four internationally distributed D-Wave annealing QPUs, demonstrating that despite architectural and parametric variations, these QPUs can be programmed to generate and validate each other’s hashes in a way that allowed stable blockchains to persist over thousands of broadcast blocks and hundreds of thousands of hash validations.
@@
For this demonstration, we used the critical dynamics of 3D and biclique spin glass at a fixed size as the computational problem for PoQ. This was motivated by its role in the first experimental demonstration of quantum supremacy in a real-world problem [1] that is not based on boson or random-circuit sampling. However, this choice does not constrain the quantum problem that can be used for PoQ. By choosing a sufficiently complex ensemble of experiments and sampling it cryptographically, experimental outcomes can be made weakly correlated. Strategies, such as randomizing problem topology, dimensionality, qubit count, and annealing time, can prevent classical learning algorithms from predicting expectation values based on large datasets [50]. Additionally, hash algorithms leveraging entanglement witnesses and shadow tomography can further enhance the quantum nature of the hashing process. We note that while supremacy is desirable for PoQ, it may not be strictly necessary as long as classical competition remains prohibitively expensive, making quantum approaches the only practicable ones. In fact, one may deliberately use smaller problem instances where the ground truth can be obtained — albeit with significant computational effort — to benchmark and calibrate QPUs.
@@
A quantum blockchain could transform existing blockchain technology by significantly mitigating the substantial electricity demands of classical proof of work (PoW) mining. Electricity costs are estimated to comprise 90% to 95% of Bitcoin’s total mining expenses [51]. In our PoQ approach, energy consumption amounts to a small fraction of quantum computation cost (which is the dominant cost in the PoQ approach). For example, energy expenses account for just 0.1% of the total computation cost on D-Wave quantum computers [52]. Replacing classical PoW with quantum PoQ could therefore reduce energy costs by a factor of 1,000. This ratio may shift as quantum computation becomes more cost-efficient over time, but substantial energy savings are expected to persist. Transition to PoQ would also relocate mining from regions with low electricity costs to countries with advanced quantum computing infrastructure. Quantum computation has the potential to undermine the security of some proof-of-work blockchains, unstructured database search (Grover’s algorithm) being the best known example [53]. Unitary evolution in the quantum-supremacy regime is at the core of our algorithm; Grover’s algorithm requires an efficient oracle to determine membership, which cannot be constructed in our case, and thus a layer of security is provided. Beyond demonstrating the first practical use of quantum supremacy in a critical domain, this work proves the feasibility of quantum computing in real-world tasks and suggests broader high-impact applications with today’s technology.
@@ Acknowledgment
We thank Brian Barch, Vladimir Vargas Calderon, Rahul Deshpande, Pau Farré, Fiona Hannington, Richard Harris, Emile Hoskinson, Trevor Lanting, Daniel Lidar, Peter Love, Chris Rich, Murray Thom and Gunnar Miller for fruitful discussions and comments on the manuscript.
@@ Appendix A Cross-Validation Limitations in Boson and Random-Circuit Sampling
To operate within a quantum blockchain, QPUs must cross-validate each other’s outcomes [27]. Unstructured distributions yielding states with close to uniform random probability, such as those obtained from random-circuit sampling, are unsuitable for this purpose: the expectation values of all foreseeable k-local operators remain close to zero, making them ineffective in the construction of witnesses. The only observable with a nontrivial expectation is $\left\langle\psi\right|\rho\left|\psi\right\rangle$, where $\left|\psi\right\rangle$ is a ground-truth wave function. For this reason the estimator commonly used to establish supremacy is the cross entropy [38]. Given a distribution allowing fair and efficient sampling in the computational basis ($P_{\text{QPU}}$, typically a QPU), and a well-resolved ground-truth distribution in the same basis ($P_{\text{GT}}$, used for verification), the cross entropy is quantified as
@@
A.1 Coarse-grained boson sampling
A proposal for a proof-of-work blockchain using coarse-grained boson sampling (CGBS) was recently proposed by Singh et al. [29]. The form of unitary evolution and measurement that is undertaken is defined by boson sampling. In order to bypass the problem of resolving state probabilities, samples are binned; the statistic to be cross validated is then the most-probable bin, which with care and sufficiently (polynomially many) samples may be reproducible given a set of devices of comparable (high) control [27, 54]. Miners’ broadcasts are synchronized and staked (a miner pays per broadcast), and results are collected subject to a classical-filtering process (mode binning) into a single ground truth. If miners submit sample sets well-matched to the ground truth they are rewarded in proportion to the number of samples broadcast (work). Useful bounds are derived as a function of sampling error and control (photon and other losses), making clear that further development of these devices is necessary for an experimental implementation.
Binning of uncorrelated nearly-equal probability states results in a concentration of probability, so that impractically many samples are required to resolve the differences amongst bins—the problems outlined in the previous section are not bypassed by binning in the case of random-circuit sampling. However, in boson sampling evidence suggests that enough structure exists in the distribution to allow a resolution of the most-probable bin with a number of samples scaling weakly with the problem size [27]. Mode-binned distributions, and k-local correlations in the modes, allow for classical reproduction. Mode-binning is proposed by Singh et al. as a prefilter, with the final proposal being based on state-binning that is robust to classical approximation. State-binning must be done very carefully to allow efficient non-spoofable cross-validation on near-term devices, for a plurality of possible unitaries (bin permutations).
The consensus mechanism based on CGBS faces several challenges including an absence of devices operating with quantum supremacy and the need to improve quality of device. The collective ground-truth consensus mechanism proposed is quite different from the Bitcoin-like scheme we demonstrate. Whereas the evaluation of work in our proposal requires that a stakeholder trust their own verifying device, in the CGBS proposal there is a need to trust an aggregated ground truth across many miners built with a requirement for strong network synchronization. Several interesting mechanisms are proposed to mitigate for weakness of the protocol, some similar to those in Section C. In particular we note that with the combination of a stake and an efficient classical filter, many obvious attacks, based on manipulation of the ground truth, become impracticable. Further development may be require to mitigate for attacks that involve recombining, resubmitting and/or buffering samplesets for enhanced reward (attacks which even when unrewarded may destabilize verification). Note that states might be manipulated in a manner preserving exact or approximate mode populations. There are also weaknesses in our blockchain arising from, for example, the redefinition of Chainwork that do not impact this proposal.
Appendix B Witnesses construction beyond the computational basis
B.1 Multi-Basis Measurement
As an illustrative example, we consider using Bell’s inequality violations for pairs of qubits as entanglement witnesses. In an appropriate experimental setup, a loophole-free violation of Bell’s inequality indicates that the experimental results cannot be explained by any local hidden-variable theory. However, this is not our objective. Instead, we aim to use the inequality as an entanglement witness. Specifically, we employ an experimentally feasible variation of Bell’s inequality, known as the Clauser-Horne-Shimony-Holt (CHSH) inequality.
B.2 CHSH Inequality
@@
B.2.1 CHSH based witnesses
Let us consider a system in which measurements can be performed in two Pauli bases, $x$ and $z$. For any two qubits, $i$ and $j$, we define a modified Bell operator as (see also Appendix B.2):
$B_{ij}=\sqrt{2}(\sigma_{i}^{x}\sigma_{j}^{x}+\sigma_{i}^{z}\sigma_{j}^{z}).$ (35)
Based on a purely statistical argument, it is shown that for any local hidden variable theory, the expectation value of the Bell operator satisfies
$|\langle B_{ij}\rangle|\leq 2\;,$ (36)
which is known as CHSH inequality. When Pauli matrices are replaced by classical spin components, we get an even smaller bound, $|\langle B_{ij}\rangle|\leq\sqrt{2}$ (see Appendix B.2). In contrast, quantum mechanics allows $|\langle B_{ij}\rangle|$ to reach values as large as $2\sqrt{2}$. This is possible when $|\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle|=|\langle\sigma_{i}^{z}\sigma_{j}^{z}\rangle|=1$, which is classically counterintuitive as it means the spins are oriented both in $x$ and $z$ directions at the same time. Quantum mechanically it is possible because of quantum contextuality, i.e., dependence of the outcome on the measurement setup. It is clear that violation of CHSH inequality is a sufficient but not necessary condition for quantum mechanics, and more explicitly for entanglement between the qubit pair. Therefore, CHSH inequality is an entanglement witness.
CHSH inequality can also hold in a multi-qubit system and therefore can be used as a basis for hash generation. Let us define $W_{\alpha}=|\langle B_{ij}\rangle|$ in Eq. (6) with a threshold value $W_{0\alpha}=2$. Here, $\alpha$ represents a pair of qubits in the multi-qubit system. The hash bit ${\cal H}_{\alpha}$ is then equal to one if the CHSH inequality for the corresponding pair is violated, and zero otherwise. It has been shown that in a multi-qubits system the CHSH inequality must satisfy a trade-off relation [56] (see Appendix B.2):
$\sum_{\alpha\in\{{\rm all\ pairs}\}}W_{\alpha}^{2}\ \leq 2N_{Q}(N_{Q}-1)$ (37)
B.3 Shadow Tomography
$\langle O\rangle\approx{1\over N}\sum_{i}{\rm Tr}(O\hat{\rho}_{i})$ (38)
where $M$ is the shadow tomography measurement channel, defined as:
${\rm Tr}(O_{\alpha}\rho)>\kappa>{\rm Tr}(O_{\alpha}\tilde{\rho})$ (42)
@@ Appendix C Weaknesses and Mitigations
Many weaknesses for PoW blockchains such as Bitcoin are known, and many mitigations have been developed [59, 7, 8, 9, 10, 29]. We focus on several attacks that might be attempted by an adversarial agent to gain reward with incomplete work. The reward might be gained deterministically, or more often probabilistically (in expectation). Neither the attacks presented, nor the mitigations, are an exhaustive list.
@@
As described in the main text, the particular proof-of-concept experiment is premised on the notion of honest network users and is susceptible to attacks. In addition to mitigations described in the main text we explore additional procedures in this appendix. We present a set of tools to protect against no-work and partial-work attacks, but optimization of the experimental ensemble, post-processing and digitalization process, essential in practice, goes beyond the scope of this publication.
@@ C.2 Partial work attacks
An attack might proceed as follows. For many nonces an attacker can estimate $\{W_{\alpha}\}$ by classical methods (e.g. by approximate simulation, or machine learning). The results are ranked by the number of (robust) threshold-satisfying bits. The best candidate might be broadcast directly. The reward from this strategy is described as in the no-work scenario, but with an enhanced (classical) work cost and enhanced probability that the bits are verified. This attack is mitigated in part by the modified block structure of Appendix C.4. In particular, spoofing enhanced experimental data (beyond the $\mathcal{N}_{\rm zeros}$ proof of work bits) is impossible given quantum supremacy [1].
@@ C.2.1 Partial Quantum Work Attack
The quantum experimental set-up also allows room for efficiencies as well, which should be understood in a full application. To minimize QPU access time (costs) a miner should optimize parameters such as number of samples and programming time. They could then more efficiently search, or filter, nonces. A miner is incentivized to prefilter nonces with cheap experiments, and then seek higher confidence (via enhanced sampling) only on candidates for broadcast. In our proof-of-concept demonstrations by contrast we fix QPU access time to 1 second in mining and validation.
@@ C.3 Multi-Block Attack
Our definitions of Chainwork mean that for any path to be viable as a strongest chain there must be strictly more verified blocks than unverified blocks. Owing to this, the criteria necessary for this attack to succeed are similar to the standard 51% attack. In effect an attacker must be able to grow the chain more quickly than alternative non-fraudulent candidates, in order to undo the Chainwork damage of their fraudulent block(s). This requires control of a significant fraction of community compute power assuming practical chain efficiency (concentrated honest mining power, up to small fluctuations).
@@ C.5 Hyperplane distribution mitigations
Figure 7: Operation of an example blockchain of 128 block broadcasts generated in the limit of many miners using projection orthogonal to $J$ with $\mathcal{N}_{zeros}=32$, and confidence-based Chainwork with $\mathcal{N}_{\rm zeros}=1$. The most recent block mined is marked in black and the strongest chain in blue. Averaging efficiency across 5 independent chains yields an efficiency estimate of 0.34 +/- 0.03 in aggrement with the bootstrapped estimates of Figure 16.
Suppose a per-nonce classical work less than the quantum experimental cost ($C_{H}$) allows for a witness estimator ${\hat{W}}$. One mitigation is to modify the distribution of thresholds or hyperplanes (to choose $W_{0}\sim{\hat{W}}$), or to choose hyperplanes orthogonal to ${\hat{W}}$. In either case the classical approximation is rendered uninformative on the signs (bits) expected from the quantum experiment. If the uncertainty in the quantum experiment is relatively small, we might still have sufficient accuracy in quantum experiments for robust validation. A visualization of this process is shown in Fig. 12.
In the case of our quench experiments, we know that nearest-neighbor correlation are correlated with $-J$, the approximation $C_{ij}\propto-J_{ij}$ describes accurately the weak coupling (fast anneal) limit. There are significant deviations from this estimate, but any hyperplane that is well correlated with $J$ can accumulate a reliable signal from weak estimators incorporating this approximation. To mitigate for this we might choose our hyperplanes to be orthogonal to $J$. Choosing hyperplanes (or thresholds $W_{0\alpha}$) to decrease reliability of classical estimators will typically enhance the uncertainty in quantum validation, thus reducing the efficiency and enhance delays in a blockchain. This effect is shown in Fig. 16 as a function of $\mathcal{N}_{zeros}$. In Fig. 7 we show an example of a chain resistant to simple classical inference based on the weak coupling limit approximation.
@@ Appendix D Experimental Implementation Details
Figure 8: Energy scales (21) of the general-access QPUs used in experiments. $\Gamma(s)$ decreases with anneal progression $s$, ${\cal J}(s)$ increases. Transverse fields are comparable on all devices, but the problem energy scale is larger in Advantage2. Advantage QPUs can be programmed with (approximately) doubling the anneal duration to simulate an Advantage2 QPU full-energy scale anneal as shown in Figure 9 [1].
Figure 9: For Advantage systems to emulate Advantage2 annealing processes, where each operates at the maximum programmable energy scales, it is necessary to anneal for longer. These factors can be obtained by a collapse of statistics such as ensemble-average energy, here using averages on 25 random disorder realizations. Factors shown that allow a matching of energy also allow near-optimal cross-validation rates, which is consistent with the unitary dynamics being controlled by critical phenomena, and thereby weakly dependent on the detailed shape of the schedule [1].
Figure 10: Operation of an example blockchain using four generally-accessible Advantage and Advantage2 QPUs with $\mathcal{N}_{\rm zeros}=64$ and 100 miners with data collected across more than 2 weeks, with 1041 total block broadcasts and using basic +/-1 Chainwork. The color scheme matches that of Figure 3. Node placement and line-thickness conventions follow Figure 4. It is apparent in the larger number of black and gray (unconfirmed) transactions, and in the increased rate of edges spanning the spiral, that this blockchain is of lower efficiency and larger delay than that presented in Figure 4. This is consistent with larger $\mathcal{N}_{\rm zeros}$ as shown in Figure 5 and 16. Where a solver became unavailable mining and verification was redistributed on available solvers; the only significant disruption of four-solver availability was a four-hour outage on April 15th owing to a severed internet cable at the Information Sciences Institute in California (Advantage_system6.4). Although the cross validation rate fluctuated through such events, this had only a small impact on the stability of the blockchain and the measured efficiency is well matched to the bootstrapped prediction of Figure 5 at 53 %.
Figure 11: Operation of an example blockchain using the Advantage2 QPU and the dimerized biclique spin-glass problems with $\mathcal{N}_{\rm zeros}=64$ and 100 miners, 1101 block broadcasts and using basic +/-1 Chainwork. Like Figure 10, the efficiency is estimated at 53%, with data collected across more than 2 weeks. Despite differences in parameterization of the cubic and biclique experiments the median cross-validation rate is comparable, consistent with the matched efficiency. The color scheme, node placement and line thicknesses match the conventions of Figures 3 and 4
@@
Figure 13: Validation rates with respect to experiments on different QPUs. Validation rates are higher using a common QPU, or a common QPU architecture (Advantage (Adv) or Advantage2 prototype (Adv2)). They are slightly higher on Advantage QPUs due the smaller size of the prototype Advantage2 prototype QPU combined with the bound of 1 second of QPU access time. On each QPU the probability of an outcome 1 is measured on each QPU from $32$ independent (with randomized embeddings) programmings, for 20 $J$ realizations and total of $20480$ projections. From these $4\times 20\times 32$ programmings we infer bit error rates for different QPU combinations. Three Advantage QPUs are used, the relatively small variability between these QPUs is not shown. (top) Hyperplane are chosen with independent and normally distributed components. (bottom) With the same hyperplanes projected into the space orthogonal to $J$ (see Section C.5).
@@
Figure 16: Statistics from 16 chains, each of length 512-1024 using the QPU-witness bootstrapping methodology (see Appendices D.1 and D.3). Hyperplanes are either i.i.d normally distributed, or are restricted to be orthogonal-to-$J$ (strengthening hashes against weak classical estimators). Resistance to spoofing increases uncertainty in validation, reducing efficiency. The Chainwork uses either the basic +/-1 definition, or the confidence-weighted definition with $\mathcal{N}_{\rm max}=1$ or $2$: confidence-weighted strongest chains are more efficient. Statistics are established via the bootstrapping method, with statistic in agreement with the full experiments shown in Figures 4, 7 and 10. (top) Mining efficiency measures the fraction of block broadcasts entering the strongest chain in the limit of many miners. (bottom) Transactions (blocks) are uncertain until agreed upon by the majority of miners. Transaction delay with 95% confidence measures the depth required for a consolidation of 95% of mining power in the limit of many miners. Stakeholders might consider blocks buried this deep within their strongest chain as immutable with 95% confidence; higher confidence implies higher delays. Data is truncated where mining efficiency drops below 0.2, as mining depth becomes strongly biased as a function of the finite blockchain size.
@@
We performed simulations of many chains with 100 miners modeling a mining pool and transactions, or taking the limit of many miners (where no miner mines more than one block). Examples are shown in Figs. 4, 10, 11 and 7. However, only the distribution of witness statistics impacts the structure and stability of the chain, assuming no simultaneous broadcasts, via mining events and stakeholder validation. Understanding this we were able to accelerate mining (Section D.2), and simulate experiments by bootstrapping (Section D.3). We first describe our main statistical measures of blockchain stability in Section D.1.